OpenStudy (anonymous):

cosx cosy= (1/2)(sin(x+y)+sin(x-y)) A. True B. False

OpenStudy (cwrw238):

RHS = (1/2)( sinx cosy + siny cosx + sinx cos y - cosx sin y) = (1/2)(2 sinx cosy) = sinx cos y